OpenStudy (anonymous):

Evaluate: 10 + 8 ÷ 2 - 4 10 5 6 14

OpenStudy (zzr0ck3r):

pmdas

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5 is the right answer

OpenStudy (zzr0ck3r):

10+(8/2)-4 = 10+4-4=10

OpenStudy (zzr0ck3r):

no its not

OpenStudy (anonymous):

10+8=18/2=9-4=5

OpenStudy (zzr0ck3r):

google it;0

OpenStudy (zzr0ck3r):

no

OpenStudy (anonymous):

ithe right answer is 10. 10+(8:2)-4= 10+4-4=10

OpenStudy (anonymous):

Evaluate: -8 + 4

OpenStudy (anonymous):

-12 -4 4 -32

OpenStudy (anonymous):

-4

OpenStudy (zzr0ck3r):

hes right this time:)

OpenStudy (zzr0ck3r):

but close this and open new one for new questions